# On vector-valued functional equations with multiple recursive terms

### v2

@@ 3 A Markovian fluid flow model with consumption
$\begin{array}[]{rl}\mathbf{Z}(r,s,\eta,v)=&\sum_{k=0}^{n}\sum_{i_{1}+\ldots+i_{N}=k}\mathbf{F}_{i_{1},\ldots,i_{M}}(r,s,\eta,v)\mathbf{L}(r,a_{i_{1},\ldots,i_{N}}(s),\eta,v)\vskip 5.69054pt\\ &+\sum_{i_{1}+\ldots+i_{N}=n+1}\mathbf{F}_{i_{1},\ldots,i_{N}}(r,s,\eta,v)\mathbf{Z}(r,a_{i_{1},\ldots,i_{N}}(s),\eta,v),\end{array}$ (30)
where $a_{i_{1},\ldots,i_{N}}(s)=a_{1}^{i_{1}}(a_{2}^{i_{2}}(\ldots(a_{N}^{i_{N}}(s))\ldots))$ and $a_{i}^{n}(s)$ denotes the $n$th iterate of $a_{i}(s)=a_{i}s$, with $a_{0,\ldots,0}(s)=s$, and the functions $F_{i_{1},\ldots,i_{N}}(r,s,\eta,v)$ are recursively defined by
$\mathbf{F}_{i_{1},\ldots,i_{N}}(r,s,\eta)=r\sum_{k=1}^{N}\mathbf{F}_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(r,s,\eta)\tilde{P}^{(k)}\mathbf{G}(a_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s),\eta),$
@@
$\mathbf{Z}(r,s,\eta,v)=\sum_{k=0}^{\infty}\sum_{i_{1}+\ldots+i_{N}=k}\mathbf{F}_{i_{1},\ldots,i_{M}}(r,s,\eta,v)\mathbf{L}(r,a_{i_{1},\ldots,i_{N}}(s),\eta,v).$ (31)
@@
4 The Markov dependent case
In the following, we cope with generalizing the work in [20, Section 2], by further assuming that the autoregressive parameter depends also on the state of a background Markov chain. In particular, we focus on the limiting counterpart of a special case of the model considered in Section 2; see also Remark 6.
Consider a FIFO single-server queue, and let $T_{n}$ $n$th arrival to the system with $T_{1}=0$. Define also $A_{n}=T_{n}-T_{n-1}$, $n=2,3,\ldots$, i.e., is the time between the $n$th and (n1)(n-1)th arrival. Let $S_{n}$ be the service time of the $n$th arrival, $n\geq 1$. We assume that the inter-arrival and service times are regulated by an irreducible discrete-time Markov chain $\{Y_{n},n\geq 0\}$ with state space $E=\{1,2,...,N\}$ and transition probability matrix $P:=(p_{i,j})_{i,j\in E}$. Let $\tilde{\pi}:=(\pi_{1},\ldots,\pi_{N})^{T}$ be the stationary distribution of $\{Y_{n};n\geq 0\}$. Let $W_{n}$ the workload in the system just before the $n$th customer arrival. Such an arrival adds $S_{n}$ work but makes obsolete a fixed fraction $1-a_{i}$ of the work that is already present in the system, given that $Y_{n}=i\in E$. Denote also by $Z_{i}^{n}(s):=E(e^{-sW_{n}}1_{\{Y_{n}=i\}})$, $Re(s)\geq 0$, $i=1,\ldots,N$, $n\geq 0$, and assuming the limit exists, define $Z_{i}(s)=\lim_{n\to\infty}Z_{i}^{n}(s)$, $i=1,\ldots,N$. Let also $\tilde{Z}(s)=(Z_{1}(s),\ldots,Z_{N}(s))^{T}$.
Remark 6
Note that following the notation in the previous subsection, we have $A_{i,j}(s):=E(e^{-sA_{n}}1(Y_{n}=j)|Y_{n-1}=i)=p_{i,j}\frac{\lambda_{j}}{\lambda_{j}+s}$.
Remark 7
can be handled at a cost of a more complicated expression.
Theorem 8
The transforms $Z_{j}(s)$, $j=1,\ldots,N$ satisfy the system
Proof. From the recursion $W_{n+1}=[a_{i}W_{n}+B_{n}-J_{n}]^{+}$ (given $Y_{n}=i$) we obtain the following equation for the transforms $Z_{j}^{n+1}(s)$, $j=1,\ldots,N$:
$\begin{array}[]{rl}Z_{j}^{n+1}(s)=&E(e^{-sW_{n+1}}1_{\{Y_{n+1}=j\}})=\sum_{i=1}^{N}P(Y_{n}=i)E(e^{-sW_{n+1}}1_{\{Y_{n+1}=j\}}|Y_{n}=i)\vskip 5.69054pt\\ =&\sum_{i=1}^{N}P(Y_{n}=i)E(e^{-s[a_{i}W_{n}+B_{n}-J_{n}]^{+}}1_{\{Y_{n+1}=j\}}|Y_{n}=i)\vskip 5.69054pt\\ =&\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}E(e^{-s[a_{i}W_{n}+B_{n}-J_{n}]^{+}}|Y_{n+1}=j,Y_{n}=i)\vskip 5.69054pt\\ =&\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}\left[E\left(\int_{0}^{a_{i}W_{n}+B_{n}}e^{-s(a_{i}W_{n}+B_{n}-y)}\lambda_{j}e^{-\lambda_{j}y}dy|Y_{n}=i\right)+E\left(\int_{a_{i}W_{n}+B_{n}}^{\infty}\lambda_{j}e^{-\lambda_{j}y}dy|Y_{n}=i\right)\right]\vskip 5.69054pt\\ =&\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}E\left(\lambda_{j}e^{-s(a_{i}W_{n}+B_{n})}\int_{0}^{a_{i}W_{n}+B_{n}}e^{-(\lambda_{j}-s)y}dy+\int_{a_{i}W_{n}+B_{n}}^{\infty}\lambda_{j}e^{-\lambda_{j}y}dy|Y_{n}=i\right)\vskip 5.69054pt\\ =&\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}E\left(\frac{\lambda_{j}}{\lambda_{j}-s}e^{-s(a_{i}W_{n}+B_{n})}(1-e^{-(\lambda_{j}-s)(a_{i}W_{n}+B_{n})})+e^{-\lambda_{j}(a_{i}W_{n}+B_{n})}|Y_{n}=i\right)\vskip 5.69054pt\\ =&\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}E\left(\frac{\lambda_{j}e^{-s(a_{i}W_{n}+B_{n})}-se^{-\lambda_{j}(a_{i}W_{n}+B_{n})}}{\lambda_{j}-s}|Y_{n}=i\right)\vskip 5.69054pt\\ =&\sum_{i=1}^{N}p_{i,j}\left[\frac{\lambda_{j}}{\lambda_{j}-s}Z_{i}^{n}(a_{i}s)\beta_{i}^{*}(s)-\frac{s}{\lambda_{j}-s}Z_{i}^{n}(a_{i}\lambda_{j})\beta_{i}^{*}(\lambda_{j})\right]\vskip 5.69054pt\\ =&\frac{\lambda_{j}}{\lambda_{j}-s}\sum_{i=1}^{N}p_{i,j}Z_{i}^{n}(a_{i}s)\beta_{i}^{*}(s)-\frac{s}{\lambda_{j}-s}\sum_{i=1}^{N}p_{i,j}Z_{i}^{n}(a_{i}\lambda_{j})\beta_{i}^{*}(\lambda_{j}).\end{array}$
Remark 9
It is easy to realize that (34) refers to the matrix analogue of equation (2) in [2]. Moreover, $\tilde{V}(0)=\mathbf{0}$, so that $\mathbf{1}\tilde{V}(0)=0$ and $H(0)=P^{T}$, so that by denoting by $H_{j}(0)$ the $j$th column of matrix $H(0)$, we will have $\mathbf{1}H_{j}(0)=\sum_{k=1}^{N}p_{j,k}=1$, $j=1,\ldots,N$, where $\mathbf{1}$ the $1\times N$ row vector of ones. So a matrix analogue of [2, Theorem 2] should give the solution to matrix functional equations of the form as given in (34).
Iterating $n$ times (34) yields,
$\tilde{Z}(s)=\sum_{k=0}^{n}\sum_{i_{1}+\ldots+i_{N}=k}F_{i_{1},\ldots,i_{M}}(s)V(a_{i_{1},\ldots,i_{N}}(s))+\sum_{i_{1}+\ldots+i_{N}=n+1}F_{i_{1},\ldots,i_{N}}(s)\tilde{Z}(a_{i_{1},\ldots,i_{N}}(s)),$ (35)
where $a_{i_{1},\ldots,i_{N}}(s)=a_{1}^{i_{1}}(a_{2}^{i_{2}}(\ldots(a_{N}^{i_{N}}(s))\ldots))$ and $a_{i}^{n}(s)$ denotes the $n$th iterate of $a_{i}(s)=a_{i}s$, with $a_{0,\ldots,0}(s)=s$, and the functions $F_{i_{1},\ldots,i_{N}}(s)$ are recursively defined by
$F_{i_{1},\ldots,i_{N}}(s)=\sum_{k=1}^{N}F_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s)H(a_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s))\tilde{P}^{(k)},$
Note that as $n\to\infty$, $Z_{j}(a_{i_{1},\ldots,i_{N}}(s))\to Z_{j}(0)=E(1_{\{Z=j\}})=P(Z=j)=\pi_{j}$, so that $\tilde{Z}(0)=\tilde{\pi}$. Alternatively, by letting $s=0$ in (34) and having in mind that $\tilde{V}(0)=\mathbf{0}$, and $H(0)=P^{T}$, (34) yields $\tilde{Z}(0)=P^{T}\tilde{Z}(0)$, thus, $\tilde{Z}(0)=\tilde{\pi}$.
Note that for $i_{1}+\ldots+i_{N}=n$, $|a_{i_{1},\ldots,i_{N}}(s)|\leq\kappa^{n}|s|$, where $\kappa=max(a_{1},\ldots,a_{N})$. Then,
[…]
where $j^{*}\in E$ is such that $|Z_{j^{*}}(a_{i_{1},\ldots,i_{N}}(s))-\pi_{j^{*}}|\geq|Z_{j}(a_{i_{1},\ldots,i_{N}}(s))-\pi_{j}|$, $\forall j\in E$.
Rewrite (35) as follows
[…]
Note that the third term in the right hand side of (36) converges to zero as $n\to\infty$. Remind that $H(0)=P^{T}$. Thus, since for $i_{1}+\ldots+i_{N}=n$, $|a_{i_{1},\ldots,i_{N}}(s)|\leq\kappa^{n}|s|$, we have that $H(a_{i_{1},\ldots,i_{N}}(s))$ with $i_{1}+\ldots+i_{N}=n$ is close to $H(0)=P^{T}$. In other words, $H(a_{i_{1},\ldots,i_{N}}(s))-P^{T}||_{1}<\epsilon<\kappa^{-1}-1$, where $A||_{1}=max_{1\leq j\leq N}(\sum_{i=1}^{N}|a_{i,j}|)$ (the maximum absolute column sum). So there is a constant $C$ such that for $i_{1}+\ldots+i_{N}=n$, $F_{i_{1},\ldots,i_{N}}(s)||_{1}\leq\binom{i_{1}+\ldots+i_{N}}{i_{1},\ldots,i_{N}}C(1+\epsilon)^{n}$. Therefore, each element in the second term in (36) is bounded by $C(1+\epsilon)^{n+1}\kappa^{n+1}|s|E(W1_{\{Y=j^{*}\}})$, which tends to zero as $n\to\infty$. The following theorem gives the main result.
Theorem 10
For $\tilde{V}(0)=\mathbf{0}$,
[…]
It remains to obtain the vector $\tilde{v}$. This is given by the following proposition.
Proposition 11
[…]
Remark 12
It would be also of high importance to consider recursions that result in the following matrix functional equations:
$\tilde{Z}(s)=H(s)\sum_{i=1}^{N}Q^{(i)}\tilde{Z}(\zeta_{i}(s))+\tilde{V}(s).$
In our case $\zeta_{i}(s)=a_{i}s$, $i=1,\ldots,N$. It seems that for general $\zeta_{i}(s)$, the analysis can be handled similarly when we ensure that they are contractions on $\{s\in\mathbb{C}:Re(s)\geq 0\}$.
4.1 The case where $A_{n+1}$, $S_{n}$ are conditionally dependent based on the FGM copula
Our aim is to obtain $Z_{j}(s;\Theta)=E(e^{-sW_{n}}1_{\{Y_{n}=j\}};\Theta)$, where $\Theta:=(\theta_{i,j})_{i,j=1,\ldots,N}$.
Theorem 13
In matrix terms,
where now,
The proof of Theorem 13 is similar to the one in Theorem 8 and further details are omitted. Following similar arguments as above, we have the following result.
Theorem 14
For $\tilde{V}(0;\Theta)=\mathbf{0}$,
[…]
$F_{i_{1},\ldots,i_{N}}(s;\Theta)=\sum_{k=1}^{N}F_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s;\Theta)U(a_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s);\Theta)\tilde{P}^{(k)},$
Proposition 15
The vectors $\tilde{v}^{(k)}$, $k=1,2,$ are given as the unique solution of the following system of equations for $j=1,\ldots,N,$:
where, for $m=1,2,$
[…]
Remark 16
Note that for $\Theta=O$, i.e., the independent copula with $\theta_{i,j}=0$, $i,j=1,\ldots,N,$ Theorem 14 is reduced to Theorem 10.
5 A modulated $D_{N}/G/1$ shot-noise queue
We now focus on the workload at arrival instants of a modulated $D_{N}/G/1$ shot-noise queue, that refers to a single server queue where the server’s speed is workload proportional, i.e., when the workload is $x$, the server’s speed equals $rx$; see [13] for a recent survey on shot-noise queueing systems, as well as [20, Section 6] that focused on the case where $N=1$. Our system operates as follows: Assume that the interarrival times $A_{1},A_{2},\ldots$ are such that $E(e^{-sA_{n}}1_{\{Y_{n-1}=i\}})=e^{-st_{i}}$, $i=1,\ldots,N$. There is a single server, and service requirements of successive customers $S_{1},S_{2},\ldots$ are i.i.d. random variables. We assume that just before the arrival of the $n$th customer, additional amount of work equal to $C_{n}$ is added. This can be explained as random noise caused by the arrival, and may be positive or negative. Let
$\begin{array}[]{rl}Z_{j}^{n+1}(s)=&E\left(e^{-sW_{n+1}}1_{\{Y_{n+1}=j\}}\right)=\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}E\left(e^{-sW_{n+1}}|Y_{n+1}=j,Y_{n}=i\right)\vskip 5.69054pt\\ =&\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}\left[pE\left(e^{-s(e^{-rA_{n+1}}(W_{n}+S_{n})+C^{+}_{n+1})}|Y_{n+1}=j,Y_{n}=i\right)\right.\vskip 5.69054pt\\ &\left.+qE\left(e^{-s[e^{-rA_{n+1}}(W_{n}+S_{n})-C^{-}_{n+1}]^{+}}|Y_{n+1}=j,Y_{n}=i\right)\right]\vskip 5.69054pt\\ =&\sum_{i=1}^{N}p_{i,j}\left[pc_{j}^{*}(-s)\beta_{i}^{*}(se^{-rt_{i}})Z_{i}^{n}(se^{-rt_{i}})\right.\vskip 5.69054pt\\ &\left.+qP(Y_{n}=i)E\left(\int_{y=0}^{e^{-rA_{n+1}(W_{n}+S_{n})}}e^{-s(e^{-rA_{n+1}}(W_{n}+S_{n})-y)}\nu_{j}e^{-\nu_{j}y}dy\right.\right.\vskip 5.69054pt\\ &\left.\left.+\int_{y=e^{-rA_{n+1}}(W_{n}+S_{n})}^{\infty}\nu_{j}e^{-\nu_{j}y}dy|Y_{n}=i\right)\right]\vskip 5.69054pt\\ =&\sum_{i=1}^{N}p_{i,j}\left[pc_{j}^{*}(s)\beta_{i}^{*}(se^{-rt_{i}})Z_{i}^{n}(se^{-rt_{i}})\right.\vskip 5.69054pt\\ &\left.+qP(Y_{n}=i)E\left(\frac{\nu_{j}}{\nu_{j}-s}e^{-se^{-rA_{n+1}}(W_{n}+S_{n})}-\frac{s}{\nu_{j}-s}e^{\nu_{j}e^{-rA_{n+1}}(W_{n}+S_{n})}|Y_{n}=i\right)\right]\vskip 5.69054pt\\ =&\left(pc_{j}^{*}(s)+q\frac{\nu_{j}}{\nu_{j}-s}\right)\sum_{i=1}^{N}p_{i,j}\beta_{i}^{*}(se^{-rt_{i}})Z_{i}^{n}(se^{-rt_{i}})-\frac{sq}{\nu_{j}-s}\sum_{i=1}^{N}p_{i,j}\beta_{i}^{*}(\nu_{j}e^{-rt_{i}})Z_{i}^{n}(\nu_{j}e^{-rt_{i}}).\end{array}$
Letting $n\to\infty$, so that $Z_{j}^{n}(s)\to Z_{j}(s)$ we have
In matrix notation, (45) is written as
where $a_{i}=e^{-rt_{i}}$, $i=1,\ldots,N$, $H(s):=\tilde{C}(s)P^{T}$. Moreover, the $n$th iterative of $\zeta_{i}(s):=a_{i}s$, i.e., $\zeta_{i}^{(n)}(s)=\zeta_{i}(\zeta_{i}(\ldots\zeta_{i}(s)\ldots))=se^{-rnt_{i}}\to 0$ as $n\to\infty$. Moreover, (47) is slightly different from (34), since the mappings $\zeta_{i}(s)$, $i=1,\ldots,N,$ are inside the summation. However, $H(e^{-rt_{i}m}s)\to P^{T}$ as $m\to\infty$, $i=1,\ldots,N$, $\tilde{Q}(0)=\mathbf{0}$. Thus, $\tilde{Z}(s)$ can be given similarly as in Theorem 10. Note that it remains to obtain the values of the vector $\tilde{r}$. This task can be accomplished using steps similar to those in Proposition 11, so further details are omitted.
Remark 17
Note that in this section we assumed that given $Y_{n}$, $Y_{n+1}$, $C_{n+1}$ and $S_{n}$ are conditionally independent. The analysis can be further applied in case we consider additional dependency, e.g., by assuming that $C_{n+1}$ and $S_{n}$ are dependent based on the FGM copula as in subsection 4.1, or have a (semi-)linear depend structure, e.g., $C_{n+1}=aS_{n}+J_{n}$, where $J_{n}$ independent random variable from $S_{n}$.
6 A modulated Markovian queue where service time depends on the waiting time
Using similar arguments as above:
$\begin{array}[]{rl}Z_{j}^{n+1}(s)=&E\left(e^{-sW_{n+1}}1_{\{Y_{n+1}=j\}}\right)=\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}E\left(e^{-s[W_{n}+[S_{n}-cW_{n}]^{+}-A_{n+1}]^{+}}|Y_{n+1}=j,Y_{n}=i\right)\vskip 5.69054pt\\ =&\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}\left[E\left(\int_{0}^{W_{n}+[S_{n}-cW_{n}]^{+}}e^{-s(W_{n}+[S_{n}-cW_{n}]^{+}-y)}\lambda_{j}e^{-\lambda_{j}y}dy|Y_{n}=i\right)\right.\vskip 5.69054pt\\ &\left.+E\left(\int_{W_{n}+[S_{n}-cW_{n}]^{+}}^{\infty}\lambda_{j}e^{-\lambda_{j}y}dy|Y_{n}=i\right)\right]\vskip 5.69054pt\\ =&\sum_{i=1}^{N}P(Y_{n}=i)p_{i,j}E\left(\frac{\lambda_{j}e^{-s(W_{n}+[S_{n}-cW_{n}]^{+})}-se^{-\lambda_{j}(W_{n}+[S_{n}-cW_{n}]^{+})}}{\lambda_{j}-s}|Y_{n}=i\right)\vskip 5.69054pt\\ =&\frac{\lambda_{j}}{\lambda_{j}-s}\sum_{i=1}^{N}p_{i,j}\left(Z_{i}^{n}(s)-\frac{s}{\mu_{i}+s}Z_{i}^{n}(s+\mu_{i}c)\right)-\frac{s}{\lambda_{j}-s}\sum_{i=1}^{N}p_{i,j}\left(Z_{i}^{n}(\lambda_{j})-\frac{\lambda_{j}}{\mu_{i}+\lambda_{j}}Z_{i}^{n}(\lambda_{j}+\mu_{i}c)\right).\end{array}$
As $n\to\infty$, $Z^{n}_{i}(s)\to Z_{i}(s)$, we have,
or equivalently,
Lemma 18
Therefore, for each $i$, the real part of $\gamma_{i}$ is positive.
Let,
$\zeta_{s}:=\{s:Re(s)\geq 0,Det(D(s))\neq 0\}.$
Then, for $s\in\zeta_{s}$,
where $A(s):=sD^{-1}(s)$, $G(s):=D^{-1}(s)\Lambda P^{T}$. Iterating (50) and having in mind that $\tilde{Z}(s)\to\mathbf{0}$ as $s\to\infty$,
and
7 A modulated ASIP tandem network with consumption
@@
Consider a queueing model consisting of two stations, say $Q_{1},Q_{2}$ in series. We assume that only $Q_{1}$ receives an external input, which is a Lévy subordinator process $X=\{X(t);t\geq 0\}$. Given the state of the background process $J(t)$, that takes values in $E=\{1,2,\ldots,N\}$, the Laplace exponent of $\{X(t);t\geq 0\}$ is $\phi_{i}(.)$, with $\phi_{i}(0)=0$, i.e., That is, $E(e^{-sX(t)}1_{\{J(t)=i\}})=e^{-\phi_{i}(s)t}$ for $s,t\geq 0$. The process $\{J(t);t\geq 0\}$ jumps to any state $j\in E$ with probability $p_{i,j}$, given that $J(t)=i$. In state $i\in E$, $J(t)$ remains for an exponentially distributed time interval with rate $q_{i}$. Given that $J(t)=i$, $Q_{k}$ has a gate that is closed except for infinitesimally short gate openings that occur at independent $exp(\mu_{k,i})$ intervals, $k=1,2$. At a gate opening of $Q_{1}$, its content instantaneously moves to $Q_{2}$; at a gate opening of $Q_{2}$, its content leaves the system. At independent $exp(\tau_{k,i})$ intervals, a fraction $a_{k,i}$ of the content of station $Q_{k}$ is instantaneously removed from the system, $k=1,2$. Let $(Z_{1}(t),Z_{2}(t))$ be the buffer contents of $(Q_{1},Q_{2})$ at time $t\geq 0$, with $Z_{1}(0)=0=Z_{2}(0)$, and with LST $f_{i}(t,s_{1},s_{2})=E(e^{-s_{1}Z_{1}(t)-s_{2}Z_{2}(t)}1_{\{J(t)=i\}})$, $i\in E$. Then, for $h\to 0$ we have,

### v3

@@ 3.6 A Markovian fluid flow model with consumption
$\begin{array}[]{rl}Z(r,s,\eta,v)=&\sum_{k=0}^{n}\sum_{i_{1}+\ldots+i_{N}=k}F_{i_{1},\ldots,i_{M}}(r,s,\eta,v)L(r,\alpha_{i_{1},\ldots,i_{N}}(s),\eta,v)\vskip 5.69054pt\\ &+\sum_{i_{1}+\ldots+i_{N}=n+1}F_{i_{1},\ldots,i_{N}}(r,s,\eta,v)Z(r,\alpha_{i_{1},\ldots,i_{N}}(s),\eta,v),\end{array}$ (68)
where $\alpha_{i_{1},\ldots,i_{N}}(s)=\alpha_{1}^{i_{1}}(\alpha_{2}^{i_{2}}(\ldots(\alpha_{N}^{i_{N}}(s))\ldots))$ and $\alpha_{i}^{n}(s)$ denotes the $n$th iterate of $\alpha_{i}(s)=a_{i}s$, with $a_{0,\ldots,0}(s)=s$, and the functions $F_{i_{1},\ldots,i_{N}}(r,s,\eta,v)$ are recursively defined by
$F_{i_{1},\ldots,i_{N}}(r,s,\eta)=r\sum_{k=1}^{N}F_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(r,s,\eta)\tilde{P}^{(k)}G(\alpha_{i_{1},\ldots,i_{k}-1,\ldots,i_{N}}(s),\eta),$
@@
$Z(r,s,\eta,v)=\sum_{k=0}^{\infty}\sum_{i_{1}+\ldots+i_{N}=k}F_{i_{1},\ldots,i_{M}}(r,s,\eta,v)L(r,\alpha_{i_{1},\ldots,i_{N}}(s),\eta,v).$ (69)
@@
4 Multidimensional extensions
In the following, we cope with some application examples, the analysis of which lead to multidimensional versions of (1). In particular, we first investigated a modulated ASIP tandem network of two queues with consumption, and then, we focus on a vector-valued reflected autoregressive process. Although the derived functional equations are (more or less) different, the solution machinery is similar to that in (1). In particular, the analysis of an instance of the modulated ASIP tandem network with consumption results in a functional equation of the form (4), where $\alpha_{i}(t):=b_{i}(t)$, $b_{i}\in(0,1)$, $i=1,\ldots,N$, in which we have two coupled recursions; see subsection (4.1). Its solution is given by Theorem 3. The analysis of the VAR(1) model results in a different vector-valued functional equation, which is far more interesting. However, the solution approach of a special case has similarities with the machinery applied in Section 2; see subsection 4.2 and Theorem 32.
4.1 A modulated ASIP tandem network with consumption
@@
Consider a queueing model consisting of two stations, say $Q_{1},Q_{2}$ in series. We assume that only $Q_{1}$ receives an external input, which is a Lévy subordinator process $X=\{X(t);t\geq 0\}$. Given the state of the background process $J(t)$, that takes values in $E=\{1,2,\ldots,N\}$, the Laplace exponent of $\{X(t);t\geq 0\}$ is $\phi_{i}(.)$, with $\phi_{i}(0)=0$, i.e., that is, $\mathbb{E}(e^{-sX(t)}1_{\{J(t)=i\}})=e^{-\phi_{i}(s)t}$ for $s,t\geq 0$. The process $\{J(t);t\geq 0\}$ jumps to any state $j\in E$ with probability $p_{i,j}$, given that $J(t)=i$. In state $i\in E$, $J(t)$ remains for an exponentially distributed time interval with rate $q_{i}$. Given that $J(t)=i$, $Q_{k}$ has a gate that is closed except for infinitesimally short gate openings that occur at independent $exp(\mu_{k,i})$ intervals, $k=1,2$. At a gate opening of $Q_{1}$, its content instantaneously moves to $Q_{2}$; at a gate opening of $Q_{2}$, its content leaves the system. At independent $exp(\tau_{k,i})$ intervals, a fraction $a_{k,i}$ of the content of station $Q_{k}$ is instantaneously removed from the system, $k=1,2$. Let $(Z_{1}(t),Z_{2}(t))$ be the buffer contents of $(Q_{1},Q_{2})$ at time $t\geq 0$, with $Z_{1}(0)=0=Z_{2}(0)$, and with LST $f_{i}(t,s_{1},s_{2})=\mathbb{E}(e^{-s_{1}Z_{1}(t)-s_{2}Z_{2}(t)}1_{\{J(t)=i\}})$, $i\in E$. Then, for $h\to 0$ we have,